Pollution. Determine the amount of pollution after the tax.
D: P = 2,400 - 2Q
S: P = 400+ 2Q
Poll.: P = 80
1
Expert's answer
2016-05-21T12:41:02-0400
If demand and supply functions for pollution are:
D: P = 1,400 - 2Q
S: P = 200 + 2Q
In equilibrium Qd = Qs, so:
1400 - 2Q = 200 + 2Q,
4Q = 1200,
Q = 300 units.
P = 1400 - 2*300 = $800.
If the price of pollution is P = $80 and Q = 300 units, then the social welfare loss before the tax is SWL = P*Q = 80*300 = $24,000.